Assume the following: Current Assets = $900; Net Fixed Assets = $2,850; Net working capital = $300; Long-term Debt = $1,500. What is Owners’ Equity?
Answer #1

Net working capital=current assets-current liabilities

current liabilities=(900-300)=$600

Total debt=current liabilities+long term debt

=600+1500=$2100

Total assets=current assets+net fixed assets

=900+2850=$3750

Total assets=Total liabilities+Total equity

Total equity=(3750-2100)=$1650.